OpenStudy (micahm):

Evaluate 24-x when x=6. Write the answer as a decimal. Round to the nearest hundredth if necessary. how would you solve that?

ganeshie8 (ganeshie8):

plugin x = 6 and simplify

ganeshie8 (ganeshie8):

\[\large 24-6\]

OpenStudy (micahm):

so i subtract that 24-6

ganeshie8 (ganeshie8):

Exactly!

OpenStudy (micahm):

but how did you get 24

OpenStudy (micahm):

the name of this lesson i am in is Rationals, Irrationals, and Radicals

ganeshie8 (ganeshie8):

Okay, could you take a screenshot of the question and attach if psble ?

OpenStudy (micahm):

how do you take a screen shot on a PC

ganeshie8 (ganeshie8):

press that, then open paint and paste it

OpenStudy (micahm):

i press what it told me but nothing happened

OpenStudy (amistre64):

or use the draw feature ....

OpenStudy (micahm):

\[\sqrt{24-x} when x=6 \]

OpenStudy (micahm):

Write the answer as a decimal. Round to the nearest hundredth if necessary.

ganeshie8 (ganeshie8):

Oh so you have that radical in the original expression

ganeshie8 (ganeshie8):

still the process is same - simply plug in x = 6 and evaluate

ganeshie8 (ganeshie8):

\[\large \sqrt{24-6}\]

OpenStudy (micahm):

A. 1.41 B. 3 C.4.24 D. 9

OpenStudy (micahm):

these are what i have to choose from A. 1.41 B. 3 C.4.24 D. 9

OpenStudy (amistre64):

which 2 perfect squares does this come between?

OpenStudy (micahm):

as in

OpenStudy (amistre64):

as in, the squart root of 24-6 falls between which 2 perfect squares

OpenStudy (micahm):

i have to subtract right and if i do that it gives me 18

OpenStudy (amistre64):

good now 18 isnt a perfect square, its an integer whose square root is not an integer but it falls between 2 perfect squares can you make a small list of perfect suqares?

OpenStudy (micahm):

so 24-6 has two perfect square to let you know i need an example i am not that great with math but an example will help me

OpenStudy (amistre64):

24-6 = 18 and we want to know where 18 fits in a list of perfect squares ... 1^2 = 1 2^2 = 4 3^2 = 9 4^2 = 16 (**)^2 = 18 <------ 5^2 = 25

OpenStudy (amistre64):

18 fits between 16 and 25, so the sqrt(18) fits between 4 and 5

OpenStudy (amistre64):

only one of your options is bigger than 4, but less than 5

OpenStudy (micahm):

9 is bigger and 3 is less

OpenStudy (amistre64):

so we know the answer aint 9 or 3

OpenStudy (micahm):

1.41 is less 4.29 is between 5

OpenStudy (amistre64):

so we should go with 4.29, do you agree?

OpenStudy (micahm):

i agree cus it's less then 5 but at the same time it's not larger then 5 thank very much